joshiscool:

What is the point slope form of the line with slope −14 that passes through the point (−2, 9) ? y+9=−14(x−2) y+2=−14(x−9) y−9=−14(x+2) y−2=−14(x+9)

joshiscool:

@Allison

Allison:

I already tagged everyone on the other post D:

loveboy:

slope,m=-14 point,(x,y)=(-2,9) as we know that point slope form is represented as:- \[\large \bf y-y_1=m(x-x_1)\] so plug in the values we get \[\large \bf y-9=-14(x+2)\]

loveboy:

this leads to option C :)